# A Resolution Independent Neural Operator

### v2

@@ Abstract
The Deep operator network (DeepONet) is a powerful yet simple neural operator architecture that utilizes two deep neural networks to learn mappings between infinite-dimensional function spaces. This architecture is highly flexible, allowing the evaluation of the solution field at any location within the desired domain. However, it imposes a strict constraint on the input space, requiring all input functions to be discretized at the same locations; this limits its practical applications. In this work, we introduce a Resolution Independent Neural Operator (RINO) that provides a framework to make DeepONet resolution-independent, enabling it to handle input functions that are arbitrarily, but sufficiently finely, discretized. To this end, we propose two dictionary learning algorithms to adaptively learn a set of appropriate continuous basis functions, parameterized as implicit neural representations (INRs), from correlated signals defined on arbitrary point cloud data. These basis functions are then used to project arbitrary input function data as a point cloud onto an embedding space (i.e., a vector space of finite dimensions) with dimensionality equal to the dictionary size, which can be directly used by DeepONet without any architectural changes. In particular, we utilize sinusoidal representation networks (SIRENs) as trainable INR basis functions. The introduced dictionary learning algorithms can be used in a similar way to learn an appropriate dictionary of basis functions for the output function data. This approach can be seen as an extension of POD-DeepONet for cases where the realizations of the output functions have different discretizations, making the Proper Orthogonal Decomposition (POD) approach inapplicable. We demonstrate the robustness and applicability of RINO in handling arbitrarily (but sufficiently richly) sampled input and output functions during both training and inference through several numerical examples.
@@ 1 Introduction
Figure 1: Vanilla DeepONet and RINO in a nutshell. Vanilla DeepONet requires input functions to be sampled at a fixed set of points (A.1), while RINO allows for arbitrary discretization of the input functions (B.1). The branch network in Vanilla DeepONet operates directly on the discretized input functions (A.2), whereas in RINO, the DeepONet operates on the embeddings obtained from projecting the input signal onto a dictionary learned a priori (B.2).
@@
Modern operator learning frameworks employing deep neural networks (DNNs) have emerged as a powerful method for constructing discretization-independent emulators for PDEs that describe physical systems. Neural operators, first introduced in $2019$ with the deep operator network (DeepONet) [22], utilize two DNNs to create a surrogate model, enabling fast inference and high generalization accuracy. In this architecture one DNN, known as the branch network, encodes the input functions at fixed sensor points (see Figure 1), while another DNN, the trunk network, encodes information related to the output function at various spatiotemporal coordinates. The solution operator is defined as a linear combination of basis functions learned by the trunk network, with their coefficients learned by the branch network. The design of DeepONet is inspired by the universal approximation theorem for operators proposed by Chen & Chen [9]. The generalized theorem for DeepONet essentially replaces shallow networks used for the branch and trunk net in [9] with DNNs to gain expressivity. Since its introduction, DeepONet has been employed to address challenging problems involving complex high-dimensional dynamical systems [11, 5, 17, 39, 7]. The framework is highly versatile, allowing the two DNNs to adopt various architectures such as convolutional neural networks (CNNs) or multilayer perceptrons (MLPs), depending on the specific problem requirements.
@@
Despite its significant success, a notable limitation of the vanilla DeepONet framework is the need to discretize all input functions in the training and test sets at fixed sensor locations. This requirement can be a major bottleneck in some applications, either in cases where different computational meshes are used or where a single computational mesh is adaptively refined in time or space. In this work, we develop efficient strategies for creating unique continuous representations of input functions without constraints on their discretization, such as the number and location of sensors in the input space of both the labeled training dataset and the testing samples as shown in Figure 1, that can be consistently parameterized and applied directly to existing DeepONet architectures. This results in the new Resolution Independent Neural Operator (RINO) variant of DeepONet. Note, however, that although they do not require fixed sensor locations, these strategies of course require sufficient resolution to capture the characteristic length or time scales of the underlying function. For example, discretization must still satisfy the classical restriction on the minimal sampling rate due to the Nyquist–Shannon sampling theorem [34].
@@
To achieve discretizaiton independence, we develop a data-driven algorithm that maps an arbitrarily discretized input function realization to its corresponding embedding coordinates. We construct this embedding space by projecting these input function realizations onto a set of fixed, continuous, and fully differentiable basis functions shared across all realizations, where the projection coefficients serve as the embedding coordinates. For flexibility, these basis functions are parameterized by neural networks and are learnable, although they do not have to be and could, for example, be specified a priori. Specifically, our algorithm discovers appropriate neural network basis functions from a collection of correlated signals, e.g., input function realizations. This algorithm extends standard dictionary learning techniques from finite-dimensional vector spaces to function spaces. Dictionary learning is an unsupervised machine learning method that identifies a set of basis vectors to represent data as a sparse combination of these basis vectors. It has various applications, including sparse coding [29], image denoising [24], and face recognition [46]. Unlike some classical ML methods such as principal component analysis (PCA), dictionary learning can learn overcomplete basis sets, where the number of basis vectors exceeds the dimensionality of the input space. This may provide more flexibility for signal modeling and representation.
@@
We leverage recent advances in implicit neural representations (INRs) [25, 36] to model the dictionary basis functions. These representations are continuous, differentiable functions implicitly defined and parameterized by MLPs. Instead of the explicit parameterization of the signal of interest, where the discrete signal values are the parameters (for example, signal values at 2D grids, voxels, point clouds, and meshes), INRs represent the signal values implicitly (or indirectly) as a function parameterized by a neural network. This network $\Phi:\boldsymbol{x}\rightarrow\Phi(\boldsymbol{x})$ maps spatiotemporal coordinates $\boldsymbol{x}\in\mathbb{R}^{M}$ from an $M$-dimensional space to their corresponding $N$-dimensional signal values, $\Phi(\boldsymbol{x})\in\mathbb{R}^{N}$. INRs find extensive applications in tasks such as image generation, super-resolution, object reconstruction, and modeling complex signals [36, 37]. INRs offer several advantages. First, the functions are defined on the continuous domain of $\boldsymbol{x}$ rather than on a discrete grid, enabling the representation to adapt flexibly to various resolutions. Second, they ensure that the function is fully continuous and differentiable with respect to $\boldsymbol{x}$, facilitating the use of automatic differentiation, which is crucial for problems requiring access to the function’s gradient. Lastly, they have shown greater memory efficiency compared to grid-based representations [25], as their capacity to model fine details depends on the architecture of the model rather than the grid resolution. The application of INRs as basis functions in RINO is motivated by their resolution invariance, ability to capture fine details, and differentiability. In particular, we utilize sinusoidal representation networks (SIRENs).
@@
Connection with ReNO: The problem of interest in our study is to establish a neural operator learning method that can operate with data sampled irregularly from one realization to another while maintaining relatively consistent accuracy, particularly from the input function perspective. We assume that the function realizations are bandwidth-limited signals and that the density of sensor locations is sufficiently fine, exceeding the minimal Nyquist-Shannon sampling rate. Therefore, we do not explore potential aliasing issues related to sampling below the Nyquist-Shannon rate during training. Representation equivalent Neural Operators (ReNO) [3] were developed to investigate aliasing issues in the context of operator learning and proposes a general formalism based on frame theory to determine the conditions under which a discrete representation of an operator is equivalent to its continuous version. Since our dictionary learning algorithm identifies basis functions that are (weakly) orthogonal, they may form a frame; Bartolucci et al. [3] show that orthogonality is a sufficient condition for defining a frame. The connection between discrete and continuous representations is established by these frames (basis functions) via the synthesis operator (reconstruction operator) and its adjoint, known as the analysis operator (projection operator). Therefore, the RINO framework aligns with their frame theory, but the proposed RINO further provides an algorithm to adaptively construct appropriate basis functions (frames) to enable consistent mapping between discrete and continuous representations of signals of interest.
@@
The paper is organized as follows. In Section 2, we describe the components of our proposed framework, namely the operator learning architecture, the concept of learning a dictionary of basis functions, and the parametrization approach via the INRs. In Section 3, we present four numerical examples to test the RINO framework. For each numerical example, we conducted a comprehensive study by varying the input function, the number of sensors, and their locations. Details about the data generation process are provided in each respective example. Finally, we summarize our observations and provide concluding remarks in Section 4. An important discussion on the connection between our proposed function-space dictionary learning algorithm and classical Gappy POD (GPOD), as well as alternative possibilities using random projections to achieve vector representations of input or output function data, is presented in Appendix F.
@@ 2 Formulation
In Section 2.1, we revisit the operator learning problem statement that is the focus of this work. We utilize the DeepONet formalism to learn operators in a data-driven manner and develop a formulation that accepts a resolution-independent embedding of the input functions (e.g., spatially and/or temporally dependent source terms, initial conditions, material properties, etc.) instead of directly working with input function data in the original space. This approach eliminates the resolution dependency of the vanilla DeepONet from the perspective of the input function. At the end of this section, we present an operator learning formulation that bypasses the need for an unknown trunk during the operator learning training. This formulation extends the POD-DeepONet approach into a fully resolution-independent framework. Then, in Section 2.2, we introduce two algorithms to learn appropriate parametrized basis functions to represent input or output function data in a resolution-independent manner. Finally, in Section 2.3, we specifically introduce the parametrization of our basis functions as INRs.
@@ 2.1 Operator Learning Between Function Spaces
$\mathcal{G}\left(u\right)(y)\approx\mathcal{F}_{\theta}(u)(y)\mathrel{\stackrel{{\scriptstyle\text{DeepONet}}}{{=}}}{\mathcal{F}_{\boldsymbol{\theta}}^{\text{br}}}^{T}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bar{u}})\ {\mathcal{F}_{\boldsymbol{\theta}}^{\text{tr}}}(y)=\sum_{k=1}^{P}\text{br}_{k}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bar{u}};\boldsymbol{\theta})\ \text{tr}_{k}(y;\boldsymbol{\theta}),$ (4)
@@
One approach to mitigate this issue is to find a vector representation of the input data that is resolution-independent. If such a representation exists, we hypothesize that the classical DeepONet can be formulated in this embedding space as follows:
$\mathcal{G}\left(u\right)(y)\approx\sum_{k=1}^{P}\text{br}_{k}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\boldsymbol{\alpha}(u)};\boldsymbol{\theta})\ \text{tr}_{k}(y;\boldsymbol{\theta}),$ (5)
where $\boldsymbol{\alpha}\in\mathbb{R}^{Q}$ is an embedding of the input function $u(x)$ in a well-defined embedding space; see Figure 1(B.2) for a schematic of the architecture and Figure 2 for a schematic of the function space mappings. Note that the selection of the embedding can be defined arbitrarily by projecting the input onto a set of continuous orthogonal basis functions (e.g. orthogonal polynomials). However, we aim to identify an optimally compact basis that both improves accuracy and reduces the number of necessary parameters in the operator. In the next section, we will introduce a method to find this desired representation.
@@
2.1.1 Operator Learning with Predefined Trunk
In the vanilla DeepONet, the trunk network essentially tries to find appropriate basis functions that approximate the output functions. So, if a suitable set of bases already exists for the output function data, it can be directly used. Using this intuition, Lu et al. [23] propose the POD-DeepONet, where the trunk network is replaced by fixed, predefined Proper Orthogonal Decomposition (POD) modes associated with the output functions, resulting in only the branch network being trainable. This approach may facilitate training and improve the accuracy of the operator learning task. However, due to the use of POD, which works on fixed grids, similar discretization dependence issues arise, as with the input function process in the vanilla DeepONet. For example, if the discretization of output functions varies from realization to realization, POD cannot be directly used.
In RINO, we develop dictionary learning algorithms that adaptively learn basis functions for approximating signals defined on arbitrary point clouds. This allows us to project arbitrary signals onto these basis functions and use their corresponding coefficients as appropriate representations in the embedding space. This is the nature of the embedding of the input function in Eq. (5). If such a method exists (as discussed in the next section), it can also be applied separately to the output function data. This process would build an appropriate dictionary of basis functions for the output data, which can subsequently be used in the operator learning task as a fixed, predefined trunk (analogous to the POD) as showed in Figure 2(b).
Figure 2: Operator learning with RINO for (a) unknown trunk and (b) predefined trunk basis functions. In (a), only the input functions are embedded using a learned set of basis function and the operator is learned between the embedding space and the output directly by training both branch ($\mathcal{F}_{\boldsymbol{\theta}}^{\text{br}}$) and trunk ($\mathcal{F}_{\boldsymbol{\theta}}^{\text{br}}$) networks. In (b), is functions are learned separately for both output and input function data. Consequently, the operator learning task effectively becomes a mapping between these two constructed subspaces via the branch network $\mathcal{F}_{\boldsymbol{\theta}}^{\text{br}}$.

### v3

@@ Abstract
The Deep operator network (DeepONet) is a powerful yet simple neural operator architecture that utilizes two deep neural networks to learn mappings between infinite-dimensional function spaces. This architecture is highly flexible, allowing the evaluation of the solution field at any location within the desired domain. However, it imposes a strict constraint on the input space, requiring all input functions to be discretized at the same locations; this limits its practical applications. In this work, we introduce a general framework for operator learning from input-output data with arbitrary number and locations of sensors. This begins by introducing a resolution-independent DeepONet (RI-DeepONet), enabling it to handle input functions that are arbitrarily, but sufficiently finely, discretized. To this end, we propose two dictionary learning algorithms to adaptively learn a set of appropriate continuous basis functions, parameterized as implicit neural representations (INRs), from correlated signals defined on arbitrary point cloud data. These basis functions are then used to project arbitrary input function data as a point cloud onto an embedding space (i.e., a vector space of finite dimensions) with dimensionality equal to the dictionary size, which can be directly used by DeepONet without any architectural changes. In particular, we utilize sinusoidal representation networks (SIRENs) as trainable INR basis functions. The introduced dictionary learning algorithms are then used in a similar way to learn an appropriate dictionary of basis functions for the output function data, which defines a new neural operator architecture referred to as the Resolution Independent Neural Operator (RINO). In the RINO, the operator learning task simplifies to learning a mapping from the coefficients of input basis functions to the coefficients of output basis functions. We demonstrate the robustness and applicability of RINO in handling arbitrarily (but sufficiently richly) sampled input and output functions during both training and inference through several numerical examples.
@@ 1 Introduction
Figure 1: Vanilla DeepONet and RI-DeepONet in a nutshell. Vanilla DeepONet requires input functions to be sampled at a fixed set of points (A.1), while RI-DeepONet allows for arbitrary discretization of the input functions (B.1). The branch network in Vanilla DeepONet operates directly on the discretized input functions (A.2), whereas in RI-DeepONet, the DeepONet operates on the embeddings obtained from projecting the input signal onto a dictionary learned a priori (B.2).
@@
Modern operator learning frameworks employing deep neural networks (DNNs) have emerged as a powerful method for constructing emulators for PDEs that describe physical systems. Neural operators, first introduced in $2019$ with the deep operator network (DeepONet) [34], utilize two DNNs to create a surrogate model, enabling fast inference and high generalization accuracy. In this architecture one DNN, known as the branch network, encodes the input functions at fixed sensor points (see Figure 1), while another DNN, the trunk network, encodes information related to the output function at various spatiotemporal coordinates. The solution operator is defined as a linear combination of basis functions learned by the trunk network, with their coefficients learned by the branch network. The design of DeepONet is inspired by the universal approximation theorem for operators proposed by Chen & Chen [11]. The generalized theorem for DeepONet essentially replaces shallow networks used for the branch and trunk net in [11] with DNNs to gain expressivity. Since its introduction, DeepONet has been employed to address challenging problems involving complex high-dimensional dynamical systems [13, 7, 20, 57, 9]. The framework is highly versatile, allowing the two DNNs to adopt various architectures such as convolutional neural networks (CNNs) or multilayer perceptrons (MLPs), depending on the specific problem requirements.
@@
Despite its significant success, a notable limitation of the vanilla DeepONet framework is the need to discretize all input functions in the training and test sets at fixed sensor locations. This requirement can be a major bottleneck in some applications; for example in cases where different computational meshes are used, multi-fidelity training data are provided, or where a single computational mesh is adaptively refined in time or space. In this work, we develop efficient strategies for creating unique continuous representations of input and output functions without constraints on their discretization, such as the number and location of sensors for both the labeled training dataset and the testing samples that can be consistently parameterized and applied directly to existing DeepONet architectures (Resolution-Independent DeepONet, RI-DeepONet in Figure 1) or used to define a new neural operator in the embedding space (Figure 2). This results in the new Resolution Independent Neural Operator (RINO). Note, however, that although they do not require fixed sensor locations, these strategies of course require sufficient resolution to capture the characteristic length or time scales of the underlying function. For example, discretization must still satisfy the classical restriction on the minimal sampling rate due to the Nyquist–Shannon sampling theorem [51].
@@
To achieve resolution independence, we develop a data-driven algorithm that maps an arbitrarily discretized input function realization to its corresponding embedding coordinates. We construct this embedding space by projecting these input function realizations onto a set of fixed, continuous, and fully differentiable basis functions shared across all realizations, where the projection coefficients serve as the embedding coordinates. This approach is motivated by and rooted in classical numerical analysis concepts, such as the Stone–Weierstrass theorem [56], providing interpretability for the learned representation in terms of the identified basis functions. For flexibility, these basis functions are parameterized by neural networks and are learnable, although they do not have to be and could, for example, be specified a priori. Specifically, the proposed algorithm discovers appropriate neural network basis functions from a collection of correlated signals, i.e., input/output function realizations. This algorithm extends standard dictionary learning techniques from finite-dimensional vector spaces to function spaces. Dictionary learning is an unsupervised machine learning method that identifies a set of basis vectors to represent data as a sparse combination of these basis vectors. It has various applications, including sparse coding [42], image denoising [36], and face recognition [64]. Unlike some classical ML methods such as principal component analysis (PCA) [44, 23] or Proper Orthogonal Decomposition (POD) [4, 27], dictionary learning can learn overcomplete basis sets, where the number of basis vectors exceeds the dimensionality of the input space. This may provide more flexibility for signal modeling and representation.
@@
We leverage recent advances in implicit neural representations (INRs) [37, 53] to model the dictionary basis functions. These representations are continuous, differentiable functions implicitly defined and parameterized by MLPs. Instead of the explicit parameterization of the signal of interest, where the discrete signal values are the parameters (for example, signal values at 2D grids, voxels, point clouds, and meshes), INRs represent the signal values implicitly (or indirectly) as a function parameterized by a neural network. This network $\Phi:\boldsymbol{x}\rightarrow\Phi(\boldsymbol{x})$ maps spatiotemporal coordinates $\boldsymbol{x}\in\mathbb{R}^{M}$ from an $M$-dimensional space to their corresponding $N$-dimensional signal values, $\Phi(\boldsymbol{x})\in\mathbb{R}^{N}$. INRs find extensive applications in tasks such as image generation, super-resolution, object reconstruction, and modeling complex signals [53, 54]. INRs offer several advantages. First, the functions are defined on the continuous domain of $\boldsymbol{x}$ rather than on a discrete grid, enabling the representation to adapt flexibly to various resolutions. Second, they ensure that the function is fully continuous and differentiable with respect to $\boldsymbol{x}$, facilitating the use of automatic differentiation, which is crucial for problems requiring access to the function’s gradient. Lastly, they have shown greater memory efficiency compared to grid-based representations [37], as their capacity to model fine details depends on the architecture of the model rather than the grid resolution. The application of INRs as basis functions in RINO is motivated by their resolution invariance, ability to capture fine details, and differentiability. In particular, we utilize sinusoidal representation networks (SIRENs) [53].
@@
Connection with ReNO: The problem of interest in this study is to establish a neural operator learning method that can operate with input-output data sampled irregularly from one realization to another while maintaining relatively consistent accuracy. We assume that the function realizations are bandwidth-limited signals and that the density of sensor locations is sufficiently fine, exceeding the minimal Nyquist-Shannon sampling rate. Therefore, we do not explore potential aliasing issues related to sampling below the Nyquist-Shannon rate during training. Representation equivalent Neural Operators (ReNO) [3] were developed to investigate aliasing issues in the context of operator learning and proposes a general formalism based on frame theory to determine the conditions under which a discrete representation of an operator is equivalent to its continuous version. Since the proposed dictionary learning algorithm identifies basis functions that are (weakly) orthogonal, they may form a frame; Bartolucci et al. [3] show that orthogonality is a sufficient condition for defining a frame. The connection between discrete and continuous representations is established by these frames (basis functions) via the synthesis operator (reconstruction operator) and its adjoint, known as the analysis operator (projection operator). Therefore, the RINO framework aligns with the ReNO frame theory, but the proposed RINO further provides an algorithm to adaptively construct appropriate basis functions (frames) to enable consistent mapping between discrete and continuous representations of signals of interest.
Other Relevant Work: Perhaps the closest work to the proposed RINO is the recent work using neural fields for operator learning [50], but there are several fundamental differences. First, the neural fields approach relies on hypernetwork INRs [15], which consist of a single large INR network augmented with additional modulation parameters. These modulation parameters are assumed to be realization-dependent and serve as embedding coordinates. As a result, the neural field embedding becomes a highly nonlinear function of the INR hypernetwork basis, making the method less interpretable than the proposed RINO. In contrast, we construct orthogonal basis functions sequentially, ensuring they can linearly reconstruct the input or output field. This approach is deeply motivated by and connected to classical ideas in numerical analysis for function approximation, such as the Stone–Weierstrass theorem [56], providing a principled and interpretable framework. Each basis in the RINO is parameterized by a simple vanilla INR, which further enhances modularity and interpretability. Additionally, the neural field embedding is more prone to non-uniqueness, as multiple distinct modulations can produce the same reconstruction. This increases the likelihood that the neural field method requires more data for effective training. Moreover, the neural field dependence on network parameters may result in high-dimensional embeddings, further escalating the data requirements for both training and achieving generalization. Recently, Ingebrand et al. [25] used neural network basis functions to encode signals of interest for reinforcement learning applications, but again their algorithm does not enforce orthogonality. As a result, it may lack the uniqueness, compactness, and interpretability of the embeddings. Other works have explored adapting neural network architectures that are well-suited for handling unstructured data, such as transformers/attention mechanisms [31, 32, 24], graph neural networks [39], and generative adversarial networks [26].
A related approach that has been applied in the recent literature is to learn neural operators on dimension-reduced latent spaces. Hesthaven et al. [21] introduced POD-NN where POD is first employed to span the output function space using data snapshots, followed by the use of neural networks to predict the corresponding POD coefficients. Bhattacharya et al. [5] proposed PCA-based operator learning, where PCA serves as an autoencoder for both input and output function spaces, and a neural network is used to interpolate within the resulting latent spaces. However, due to the reliance of these methods on PCA or POD in the initial step, these approaches may not handle input or output data that is unaligned. Oommen et al. [43] introduced a framework for efficient operator learning in latent spaces by employing offline nonlinear dimension reduction through autoencoders and constructing a DeepONet on the resulting latent space, enabling operator learning for functions with sharp gradients. Building on this work, Kontolati et al. [29] conducted an extensive study of latent operator networks and demonstrated that training on latent spaces (Latent DeepONet or L-DeepONet) significantly accelerates operator learning while improving computational efficiency. Meanwhile, Seidman et al. [49] introduced NOMAD, which directly builds the neural operator on a nonlinear manifold of reduced dimension. While the proposed RINO is related to these works in the sense that it builds a neural operator on a latent space, its primary objective differs in an important way. Instead of prioritizing dimension reduction, which other latent neural operators do, the RINO prioritizes learning on a space of continuous and orthogonal basis functions (for both input and output functions). While the associated embedding in RINO typically reduces dimension (it does not have to), this continuous and orthogonal basis provides a unique representation for each input/output function that is resolution independent. This allows it to handle multi-fidelity data with arbitrary number and locations of sensor points.
@@
The paper is organized as follows. In Section 2, we describe the components of the proposed framework, namely the operator learning architecture, the concept of learning a dictionary of basis functions, and the parametrization approach via the INRs. In Section 3, we present six numerical examples to test the RINO framework. For each numerical example, we conducted a comprehensive study by varying the input function, the number of sensors, and their locations. Details about the data generation process are provided in each respective example. Finally, we summarize our observations and provide concluding remarks in Section 4. An important discussion on the connection between the proposed function-space dictionary learning algorithm and classical Gappy POD (GPOD), as well as alternative possibilities using random projections to achieve vector representations of input or output function data, is presented in Appendix F.
@@ 2 Formulation
In Section 2.1, we revisit the operator learning problem statement that is the focus of this work. We begin by utilizing the DeepONet formalism to learn operators in a data-driven manner and develop a formulation that accepts a resolution-independent embedding of the input functions (e.g., spatially and/or temporally dependent source terms, initial conditions, material properties, etc.) instead of directly working with input function data in the original space. This Resolution-Independent DeepONet (RI-DeepONet) approach eliminates the resolution dependency of the vanilla DeepONet from the perspective of the input function. We then present a more general resolution independent operator learning formulation, the RINO, that bypasses the need for an unknown trunk during the operator learning training and enables operator learning directly in the embedded function space. This formulation extends the POD-DeepONet approach into a fully resolution-independent framework. In Section 2.2, we introduce two algorithms to learn appropriate parametrized basis functions to represent input and/or output function data in a resolution-independent manner. Finally, in Section 2.3, we specifically introduce the basis function parametrization as INRs.
@@ 2.1 Operator Learning Between Function Spaces
$\mathcal{G}\left(u\right)(y)\approx\mathcal{F}_{\theta}(u)(y)\mathrel{\stackrel{{\scriptstyle\text{DeepONet}}}{{=}}}{\mathcal{F}_{\boldsymbol{\theta}}^{\text{br}}}^{T}(\bar{u})\ {\mathcal{F}_{\boldsymbol{\theta}}^{\text{tr}}}(y)=\sum_{k=1}^{P}\text{br}_{k}(\bar{u};\boldsymbol{\theta})\ \text{tr}_{k}(y;\boldsymbol{\theta}),$ (4)
@@
One approach to mitigate this issue is to find a vector representation of the input data that is resolution-independent. If such a representation exists, we hypothesize that a Resolution-Independent DeepONet (RI-DeepONet) can be formulated in this embedding space as follows:
$\mathcal{G}\left(u\right)(y)\approx\sum_{k=1}^{P}\text{br}_{k}(\boldsymbol{\alpha}(u);\boldsymbol{\theta})\ \text{tr}_{k}(y;\boldsymbol{\theta}),$ (5)
where $\boldsymbol{\alpha}\in\mathbb{R}^{Q}$ is an embedding of the input function $u(x)$ in a well-defined embedding space; see Figure 1(B.2) for a schematic of the architecture and Figure 2(a) for a schematic of the function space mappings of this RI-DeepONet. Note that the selection of the embedding can be defined arbitrarily by projecting the input onto a set of continuous orthogonal basis functions (e.g. orthogonal polynomials). However, we aim to identify an optimally compact basis that both improves accuracy and reduces the number of necessary parameters in the operator. In the next section, we will introduce a method to find this desired representation.
@@
2.1.1 Operator Learning Between Embedded Function Spaces (RINO)
In the vanilla DeepONet, the trunk network essentially tries to find appropriate basis functions that approximate the output functions. So, if a suitable set of bases already exists for the output function data, it can be directly used. Using this intuition, Lu et al. [35] propose the POD-DeepONet, where the trunk network is replaced by fixed, predefined POD modes associated with the output functions, resulting in only the branch network being trainable. This approach may facilitate training and improve the accuracy of the operator learning task. However, due to the use of POD, which works on fixed grids, similar discretization dependence issues arise, as with the input function process in the vanilla DeepONet. For example, if the discretization of output functions varies from realization to realization, POD cannot be directly used.
In the proposed RINO, we develop dictionary learning algorithms that adaptively learn basis functions to approximate signals defined on arbitrary point clouds. This allows us to project arbitrary signals onto these basis functions and use their corresponding coefficients as appropriate representations in the embedding space. This is the nature of the embedding of the input function in the RI-DeepONet from Eq. (5). But, if such a method exists (as discussed in the next section), it can also be applied separately to the output function data. This process would build an appropriate dictionary of basis functions for the output data, which can subsequently be used in the operator learning task as a fixed, predefined basis (analogous to the trunk in POD-DeepONet) as showed in Figure 2(b). The RINO is then built to approximate the operator that maps between these embedded functions spaces as illustrated by $\mathcal{F}_{\boldsymbol{\theta}}^{\text{ri}}$ in Figure 2(b).
Figure 2: Operator learning with RINO when the output basis functions used to span the output field are either (a) unknown or (b) fixed during the operator learning stage. In (a), only the input functions are embedded using a learned set of basis function and the operator is learned between the embedding space and the output directly by training both branch ($\mathcal{F}_{\boldsymbol{\theta}}^{\text{br}}$) and trunk ($\mathcal{F}_{\boldsymbol{\theta}}^{\text{br}}$) networks. In (b), is functions are learned separately for both output and input function data. Consequently, the operator learning task effectively becomes a mapping between these two constructed subspaces via a single nueral network $\mathcal{F}_{\boldsymbol{\theta}}^{\text{ri}}$.